OpenStudy (anonymous):

From the 11 male and 7 female sales representative for a insurance company, a team of 3 women and 3 men will be selected to attend the meeting. In how many ways can a team of 6 be selected in so many ways

OpenStudy (anonymous):

if you can just start me off..I want to learn the steps

OpenStudy (kropot72):

Groups of 3 of each gender are to be selected. The order of the 3 in each group does not matter. Therefore combinations are required. The number of combinations of the 11 men selected 3 at a time is: \[11C3=\frac{11!}{3!(11-3)!}=\frac{11!}{3!8!}\] Each of the combinations of men can be taken with each of the 7C3 combinations of women making the possible number of combinations in the complete team to be: \[\frac{11!}{3!8!}\times\frac{7!}{3!4!}=you\ can\ calculate\]

OpenStudy (anonymous):

goahead n solve it

OpenStudy (anonymous):

I think I need a break

OpenStudy (kropot72):

You should be able to do the final calculation, which is: \[Total\ number\ of\ ways=\frac{11\times10\times9\times7\times6\times5}{3\times2\times3\times2}=you\ can\ calculate\]